OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). 8x - y = 31 5x + 3y = 23

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

the answer is (4,1)

OpenStudy (anonymous):

5x+3(8x-31)=23 29x-93=23 29x=116 x=4 5*4+3y=23 3y=3 y=1 So we have (4,1)